OpenStudy (tate_langdon):

Can someone check my work If the sphere shown above has a radius of 3 units, then what is the approximate volume of the sphere? A. 9 cubic units B. 113.04 cubic units C. 56.52 cubic units D. 141.3 cubic units

OpenStudy (tate_langdon):

I say the answer is B

OpenStudy (mathstudent55):

Correct.

OpenStudy (tate_langdon):

Can you tell me why you think i'm correct @mathstudent55

OpenStudy (mathstudent55):

Because from my experience with your questions, you are either: 1. brilliant 2. lucky guesser 3. good at cheating or all of the above or a combination of the some of the above. Based on the above, you are usually correct.

OpenStudy (mathstudent55):

\(\Large V_{sphere} = \dfrac{4}{3} \pi r^3\) \(\Large V \approx \dfrac{4}{3} (3.14) (3~in.)^3 \) \(\Large V \approx 4(3.14) (3^2)~in.^3 \) \(\Large V \approx 113.04~in.^3 \)

OpenStudy (mathstudent55):

I'm not as smart as you are, so I must go back to basics and use the formula for the volume of a sphere.